One sample has a mean of M= 6 and a seond sample has a mean of M= 12. The two samples are combined into a single set of scores.&
A) what is the mean for the combined set if both of the original samples have n = 5 scores?&
B) what is the mean for the combined set if the first sample has n = 4 score and the second sample has n = 8? &
1
Expert's answer
2010-12-05T19:17:56-0500
A) The mean wil lbe equal to
X = (5*6+5*12)/(5+5) = 9
B) The mean of the combined set would be
X = (4*6 + 8*12) / (4+8) = 10.